OpenStudy (anonymous):

Fans & Medals!!! What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (jojo4eva):

y-4=-3(x-1)

OpenStudy (jojo4eva):

and you are using the formula y-y1=m(x-x1) in order to plug in your values